In each of the following cases, calculate the accounting break-even and the cash break-even points. Ignore any tax effects in calculating the cash break-even. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

    

Case Unit Price Unit Variable Cost Fixed Costs Depreciation
1 $ 2,830 $ 2,315 $ 8,000,000 $ 2,460,000
2 56 45 45,000 170,000
3 13 4 1,900 800
Case Accounting break-even Cash break-even
1
2
3
Answer #1

Case 1 Accounting Breakeven=(Depreciation + fixed cost)/( Sales - Variable)=(8000000+2460000)/(2830-2315) =20310.68
Cash break-even quantity =Fixed Cost/(Sales-Variable) =(8000000)/(2830-2315) =15533.98

Case 2 Accounting Breakeven=(Depreciation + fixed cost)/( Sales - Variable)=(170000+45000)/(56-45) =19545.45
Cash break-even quantity =Fixed Cost/(Sales-Variable) =(1700000)/(56-45) =154545.45

Cash 3  Accounting Breakeven=(Depreciation + fixed cost)/( Sales - Variable)=(1900+800)/(13-4) =300.00
Cash break-even quantity =Fixed Cost/(Sales-Variable) =1900/(13-4) =211.11